OpenStudy (anonymous):

anyone want some help in algebra or lower? 9th grade math or lower?

OpenStudy (anonymous):

its algebra 1

OpenStudy (anonymous):

can you help with 6th grade?

OpenStudy (anonymous):

yes i can

OpenStudy (anonymous):

okay i have a question i need help with

OpenStudy (anonymous):

need help with any more?

OpenStudy (anonymous):

Which equation can be used to find the answer? Usman paints the entire length of a fence in a day. He paints the last 12 m of the 38 m fence in the afternoon. How many meters of the fence does he paint in the morning? A. f = 38 ÷ 2 f = 19 19 m B. f – 12 = 38 f = 50 50 m C. f = 12 · 2 f = 24 24 m D. f + 12 = 38 f = 26 26 m

OpenStudy (anonymous):

is it D?

OpenStudy (anonymous):

D yes

OpenStudy (anonymous):

yes its D

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem any more

OpenStudy (anonymous):

@MissNoName i can help you even better if u fan me so then you can easily message me

OpenStudy (anonymous):

Which equation can be used to find the answer? An aquarium worker removed 78 fish from a tank. There were still 186 fish left in the tank. How many fish were in the tank before any were removed? A. f ÷ 2 =78 f = 156 156 fish B. f ÷ 2 =186 f = 372 372 fish C. f – 78 = 186 f = 264 264 fish D. f + 78 = 186 f = 108 108 fish

OpenStudy (anonymous):

c

OpenStudy (anonymous):

okay! =)

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

any more

OpenStudy (anonymous):

@MissNoName i can help you even better if u fan me so then you can easily message me

OpenStudy (anonymous):

I did! :P

OpenStudy (anonymous):

oh ok XD

OpenStudy (anonymous):

lmao

OpenStudy (anonymous):

need anymore help

OpenStudy (anonymous):

Which equation can be used to find the answer? Camilla needs $65 for a pair of hiking shoes. She already has $19 saved. How much more does she need to save? A. n + 19 = 65 n = 46 $46 B. n ÷ 2 = 19 n = 38 $38 C. n – 19 = 65 n = 84 $84 D. n ÷ 2 = 65 n = 130 $130

OpenStudy (anonymous):

how many more idc i love math

OpenStudy (anonymous):

A

OpenStudy (anonymous):

A

OpenStudy (anonymous):

there is that one, and one more

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

okay last one

OpenStudy (anonymous):

Which equation can be used to find the answer? Clarice spent $27 of her savings on clothes and has $49 of savings left. How much did she have in savings before buying the clothes? A. m + 27 = 49 m = 22 $22 B. m – 27 = 49 m = 76 $76 C. m ÷ 2 = 27 m = 54 $54 D. m ÷ 2 = 49 m = 98 $98

OpenStudy (anonymous):

b

OpenStudy (anonymous):

b

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

no problem, message me if you need any help

OpenStudy (anonymous):

mkay!